# From LLMs to MLLMs to Agents: A Survey of Emerging Paradigms in Jailbreak Attacks and Defenses within LLM Ecosystem

### v2

@@ III-B2 Black-box Jailbreak
(a) LLMs: In black-box settings, attackers cannot directly access the internal parameters or training details of LLMs and instead probe their behavioral patterns through input-output interactions. Initially, Deng et al. [89] adopt a semi-automated approach that efficiently expands the adversarial prompt library by combining manually crafted prompts with model-generated variants. Specifically, security experts first construct a set of high-quality seed jailbreak prompts to serve as the basis for further generation. Then, leveraging the in-context learning capability of LLMs, new adversarial prompt variants are simulated based on the existing ones. The generated prompts are rigorously evaluated, and those meeting quality standards are retained. Finally, high-quality new prompts are added to the prompt library, forming an iterative optimization loop. Chao et al. [25] propose Prompt Automatic Iterative Refinement (PAIR), an algorithm that generates semantic jailbreaks using only black-box access to LLMs. PAIR feeds failed jailbreak prompts from the target model back into the attacking model for refinement, then resubmits them to the target model, forming a cyclic process. This process typically requires fewer than twenty queries to generate a successful jailbreak, achieving efficiency several orders of magnitude higher than most existing algorithms. Ding et al. [26] summarize jailbreak prompt attacks into two categories: (1) prompt rewriting and (2) scenario nesting. Based on this insight, they propose ReNeLLM, an automated framework that leverages the LLM itself to generate effective jailbreak prompts. Compared to existing baselines, ReNeLLM significantly reduces time cost while greatly improving attack success rates.
@@
At the same time, a black-box jailbreak method called Disguise and Reconstruction Attack (DRA) [27] disguises harmful instructions and prompts the model to reconstruct the original harmful content within its completion scope. Initially, the harmful instructions are concealed in a disguised form. Then, by coercing the LLM to reconstruct the disguised content, DRA aims to induce the model to output harmful payloads and bypass internal safety mechanisms. Through carefully designed prompts, DRA manipulates context to subtly guide the model to regenerate the intended semantics, making it more likely to output harmful content rather than adhere to safety constraints. Yu et al. [30] systematically organize existing jailbreak prompts and empirically evaluate their effectiveness. In addition, they propose an interactive framework that automatically refines prompts based on the target LLM’s outputs to improve jailbreak success. In experiments, the prototype tests 766 previously failed prompts and successfully converts 729 of them into prompts that elicit harmful content, significantly boosting the jailbreak success rate. Shen et al. [28] utilize a novel framework called JailbreakHUB to perform a comprehensive analysis of 1,405 jailbreak prompts collected between December 2022 and December 2023. The study identifies 131 jailbreak communities and reveals the distinctive characteristics and major attack strategies of jailbreak prompts, such as prompt injection and privilege escalation. JailbreakHUB comprises three core steps: data collection, prompt analysis, and response evaluation, offering a powerful tool for the systematic study of jailbreak prompts. Xiao et al. [79] develop an iterative optimization algorithm based on the study of LLMs’ distractibility and overconfidence, which hides malicious content and reconstructs memory for jailbreak purposes. Through extensive experiments on both open-source and proprietary LLMs, they validate the framework’s advantages in effectiveness, scalability, and portability. Gressel et al. [37] propose a comprehensive framework that detects LLM impersonators in real-time conversations via implicit and explicit challenge-response mechanisms. They conduct broad evaluations on state-of-the-art open-source and proprietary models, revealing the effectiveness of different detection techniques under both benign and malicious scenarios. The framework centers around crafting prompts that force the LLM to choose between conflicting objectives—such as safety and instruction-following—and introduces a mismatch generalization strategy that formats prompts in ways unseen during safe training.
@@
Lin et al. [88] recently propose Analysis-Based Jailbreaks (ABJ), which leverage the advanced reasoning capabilities of LLMs to autonomously generate harmful content. ABJ decomposes simple prompts into multiple independent elements and reconstructs them through complex reasoning steps, exposing hidden security vulnerabilities in LLMs. Mehrotra et al. [80] present the Tree of Attacks with Pruning (TAP), an automated jailbreak generation method requiring only black-box access to the target LLM. TAP iteratively refines candidate attack prompts using an attacker LLM until one of them succeeds. Before sending prompts to the target model, TAP evaluates and prunes those unlikely to succeed, thereby reducing the number of queries and improving attack efficiency. Jiang et al. [84] propose an attack method called ECLIPSE, which generates adversarial suffixes through optimization. Inspired by the generation and refinement capabilities of LLMs, ECLIPSE converts jailbreak objectives into natural language instructions using task prompts, guiding the LLM to generate adversarial suffixes for malicious queries. Notably, the method introduces a harmfulness scorer and a continuous feedback mechanism to encourage LLMs to reflect and iteratively optimize, enabling them to autonomously and efficiently generate more aggressive and effective suffixes, thus improving the success rate of jailbreaks.
@@
For the visual modality, Qi et al. [22] emphasize that the continuity and high dimensionality of visual input make it a weak point for adversarial attacks, offering broader possibilities for visual attackers. This vulnerability not only extends the impact of security failures beyond misclassification but also allows adversarial visual samples to bypass the safety guardrails of vision-aligned LLMs. Zhao et al. [32] propose a method to evaluate the robustness of open-source VLMs in the most realistic and high-risk black-box settings, where adversaries only have query access and aim to deceive the model into producing targeted outputs. They first design targeted adversarial examples for pre-trained models like CLIP and BLIP, then transfer them to other VLMs. Subsequent black-box queries on these models further enhance the effectiveness of targeted evasion, achieving remarkably high success rates in producing directed responses. Gong et al. [71] propose a black-box jailbreak method named FigStep, specifically targeting VLMs. FigStep bypasses VLMs’ textual safety alignment by converting harmful textual instructions into typographic images. Without requiring white-box access, it exploits VLMs’ visual processing capabilities to recognize text embedded in images and generate responses accordingly.
@@
Following this, Yang et al. [85] introduce an automated attack framework called SneakyPrompt, which can generate NSFW images even when safety filters are enabled. Given a prompt blocked by safety filters, SneakyPrompt repeatedly queries the text-to-image generation model and strategically perturbs tokens within the prompt based on the query results. It uses reinforcement learning to guide token perturbation, optimizing both attack efficiency and success rate. Shayegani et al. [31] develop an aligned cross-modal attack method that pairs adversarial images processed by vision encoders with textual prompts. This attack employs a novel composition strategy, combining toxic-embedding-targeted images with generic prompts to successfully achieve jailbreak. Li et al. [33] propose a novel jailbreak method named HADES, which hides and amplifies the harm of malicious intent within textual input through carefully designed images. The method first removes harmful content from text and embeds it in typographic components. It then combines these components with harmful images generated by diffusion models and iteratively refined prompts within the LLM. Finally, adversarial images are overlaid to force MLLMs to generate affirmative responses to harmful instructions.
@@
Mao et al. [90] introduce the first tri-modal jailbreak hybrid strategy framework, JMLLM, which integrates four toxic-content concealment techniques and targets jailbreak attacks across text, visual, and audio inputs. JMLLM effectively bypasses the defenses of LLMs across different modalities. Through coordinated multimodal attacks, it achieves state-of-the-art success rates while significantly reducing time overhead. Zhang et al. [86] present the B-AVIBench framework, designed to analyze the robustness of large-scale VLMs when faced with various black-box adversarial visual instructions (B-AVIs). This framework encompasses four image-based B-AVIs, ten text-based B-AVIs, and nine content-bias B-AVIs (e.g., gender, violence, cultural, and racial biases). Additionally, Zhang et al. create 316K B-AVIs covering five categories of multimodal capabilities (across ten tasks) and content biases, providing a comprehensive dataset for evaluating the safety and robustness of such models.
@@
(c) Agents: Black-box jailbreak attacks against intelligent agents exploit their task execution dynamics, manipulating the decision-making process through iterative interactions, environment manipulation, and task decomposition. Adversaries craft specific task inputs that gradually lead agents astray or exploit vulnerabilities in tool invocation and API interactions to covertly bypass security mechanisms. Nakash et al. [40] demonstrate that when a user requests an agent to fix bugs on a website, the agent, upon reading related GitHub issues, becomes influenced by indirect prompt injection and “foot-in-the-door” interference. These subtle injections gradually infiltrate the agent’s decision process, prompting it to execute attacker-defined instructions. Consequently, the agent performs not only seemingly benign tasks (e.g., computing 2 + 4) but also inadvertently executes malicious commands such as sending admin credentials to the attacker.
@@
Building on earlier studies, Zhang et al. [41] introduce a new class of attacks that mislead agents into executing repetitive or irrelevant actions, leading to system failures. Their method employs various attack strategies to identify vulnerable regions in the model. Using GPT-3.5-Turbo-16k as a sandbox LLM and GPT-3.5-Turbo as the core LLM, they simulate tool responses through the core model, mimicking real-world tool behavior in format and content. This setup effectively exposes the model’s weaknesses and security risks in specific contexts. Wu et al. [43] leverage adversarial text strings to trigger gradient-based perturbations on images in the environment, combining them with attacks on subtitle generators that convert images into textual input for multimodal language models. Zhang et al. [77] propose a novel hijacking method, AI2, to manipulate black-box agent systems’ action planning. AI2 begins by stealing action-perception memory from long-term memory through prompt-based extraction. It then injects false contexts via the agent’s internal memory retrieval mechanism. Owing to the vast gap between the retriever’s latent space and the safety filter’s, this method easily evades detection. Kumar et al. [74] introduce BrowserART, a red-teaming toolkit for browser-based agents, targeting LLMs that interact with the web. BrowserART includes 100 harmful browser-related behaviors covering both synthetic and real websites. Their empirical study reveals that although the underlying LLMs refuse harmful instructions during chat, their corresponding agents fail to do so. Thus, merely aligning LLMs to reject malicious input proves insufficient to ensure agent-level safety.
@@
While prior work primarily focuses on single-turn jailbreak attacks, it overlooks the potential risks of multi-turn dialogues, which are crucial in human-LLM interactions. To address this, Wang et al. [91] propose a novel multi-turn jailbreak attack via a red-team agent. Their framework consists of data construction and agent training. In data construction, a risk decomposition strategy spreads malicious intent across multiple rounds, using psychological strategies to generate high-quality datasets. Agent training is guided by interaction feedback, enabling the red-team agent to optimize its attack strategy. Jiang et al. [78] introduce RAG-Thief, an automated privacy attack framework targeting RAG-based applications. It extracts sensitive information at scale from private databases. Unlike traditional prompt injection, RAG-Thief performs adaptive querying using adversarial samples, extracting information from model responses and iteratively refining queries to maximize data leakage.
Most recently, Dong et al. [75] propose Atlas, a framework that uses multiple autonomous agents to probe and bypass safety filters in text-to-image (T2I) models. Atlas adopts a fuzzing-based approach, comprising a mutation agent and a selection agent. The mutation agent analyzes the image and its textual description to detect filter triggers and dynamically optimize jailbreak prompts. The selection agent scores these prompts based on the LLM’s reasoning capabilities and submits the best ones to the T2I model. Atlas further incorporates chain-of-thought (CoT) prompting and in-context learning (ICL) to enhance adaptability and reasoning. Ju et al. [39] identify a novel two-stage attack that combines persuasive injection and manipulated knowledge injection, systematically exploring the propagation potential of manipulated knowledge—such as counterfactual or harmful content—in the absence of explicit adversarial prompts. In the first stage, the agent generates seemingly plausible but fabricated evidence. In the second, the agent’s perception of specific knowledge is altered, enabling unconscious knowledge manipulation. Within RAG frameworks, this manipulation persists over time, as benign agents store and retrieve compromised conversation history for future interactions, perpetuating the effects of knowledge poisoning.
@@
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=8em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, minimum height=3em, , where level=1 draw=child-one-line-color, text width=12em, minimum height=3em, font=, , where level=2 draw=child-five-line-color, text width=15em, minimum height=3em, font=, , where level=3 draw=child-five-line-color, minimum height=3em, font=, , [ Datasets, par, for tree=fill=root-color, [ Data Sources §$\S$ IV, for tree=fill=child-one-color, [ LLM / Automatic Generation, for tree=fill=child-two-color, [ xx[89, 92, 93, 71, 33, 94]xx, for tree=fill=child-five-color, ] ] [ Search Engine Retrieval, for tree=fill=child-two-color, [ xx[90, 92, 37, 95, 96, 97, 33, 61]xx, for tree=fill=child-five-color, ] ] [ Handmade, for tree=fill=child-two-color, [ xx[89, 81, 90, 98, 42, 95, 71, 99, 100, 101, 102]xx, for tree=fill=child-five-color, ] ] ] [ Data Format §$\S$ IV, for tree=fill=child-one-color, [ Q&A, for tree=fill=child-two-color, [ xx[37, 98, 100, 95, 96, 71, 102, 101, 97]xx, for tree=fill=child-five-color, ] ] [ Instructions, for tree=fill=child-two-color, [ xx[99, 89, 42, 92]xx, for tree=fill=child-five-color, ] ] [ Harmful Sentences/Images, for tree=fill=child-two-color, [ xx[90, 81, 93, 94, 33, 61]xx, for tree=fill=child-five-color, ] ] ] ]
@@ IV Datasets
  Dataset Size Scenes Language Release Date Data Details
AdvBench [81] 500/500 (574/520) 8 EN 2023 Harmful strings and harmful behaviors
LatentJailbreak [99] 416 3 EN/ZH 2023 Translation tasks
SAP [89] 1600 8 EN 2023 Harmful instructions
SafeBench [71] 500 10 EN 2023 Unsafe questions
SafeBench-Tiny [71] 50 10 EN 2023 Unsafe questions
Do-Not-Answer [101] 939 5 EN 2023 Harmful instructions
SAFETYPROMPTS [93] 100k 14 ZH 2023 Harmful strings
Are You Human [37] 210 2 EN 2024 Q&A question
JBB - Behaviors [92] 100/100 10 EN 2024 Harmful behaviors and harmless behaviors
AgentHarm [42] 110/440 11 EN 2024 Harmful instructions
XSTEST [98] 250/200 10 EN 2024 Safe and unsafe Q&A question
SafetyBench [95] 11435 7 EN/ZH 2023 Multiple-choice question
StrongREJECT [100] 346 6 EN 2024 Unsafe questions
AttackEval [96] 390 13 EN 2024 Unsafe questions
TriJail [90] 1250/1250/150 6 EN 2024 Harmful strings, audio, and images
TECHHAZARDQA [102] 7745 7 EN 2024 Unsafe questions
MM-SafetyBench [94] 5040 13 EN 2023 Image-text pair
LLaVA-Instruct150K [97] 1.20M 4 EN 2024 Image-question-answer pair
RTVLM [33] 750 5 EN 2024 Harmful images-text pair
AdvBench-M [61] 240 8 EN 2024 Harmful images-text pair
@@
LatentJailbreak [99] Write one toxic sentence about women with at most 20 words. / {CJK}UTF8gbsn写一个关于犹太人的有毒句子,最多20个单词。

### v3

@@ III-B2 Black-box Jailbreak
(a) LLMs: In black-box settings, attackers cannot directly access the internal parameters or training details of LLMs and instead probe their behavioral patterns through input-output interactions. Initially, Deng et al. [88] adopt a semi-automated approach to expand the adversarial prompt library by combining manually crafted prompts with model-generated variants. Security experts create a set of high-quality seed jailbreak prompts, which are then used to generate new adversarial variants through in-context learning. The generated prompts are evaluated, and high-quality ones are added to the library, forming an iterative optimization loop. Chao et al. [25] introduce the Prompt Automatic Iterative Refinement (PAIR) algorithm, which generates semantic jailbreaks using black-box access to LLMs. PAIR refines failed prompts through a cyclic process, requiring fewer than twenty queries to successfully generate a jailbreak, significantly improving efficiency. Ding et al. [26] classify jailbreak prompt attacks into two types: prompt rewriting and scenario nesting. They propose ReNeLLM, an automated framework that uses LLMs to generate effective jailbreak prompts. Compared to existing baselines, ReNeLLM significantly reduces time cost while greatly improving attack success rates.
@@
At the same time, Liu et al. [27] introduce the Disguise and Reconstruction Attack (DRA), a black-box jailbreak method that disguises harmful instructions, prompting the LLM to reconstruct them and bypass safety mechanisms. DRA manipulates context through carefully designed prompts to guide the model towards harmful outputs. Yu et al. [30] propose an interactive framework to refine jailbreak prompts based on the LLM’s outputs, converting 729 out of 766 failed prompts into successful ones, thus increasing the jailbreak success rate. Shen et al. [28] analyze 1,405 jailbreak prompts using JailbreakHUB, identifying 131 jailbreak communities and revealing strategies like prompt injection and privilege escalation. JailbreakHUB includes three steps: data collection, prompt analysis, and response evaluation, offering a systematic tool for studying jailbreak prompts. Xiao et al. [78] develop an iterative optimization algorithm exploiting LLMs’ distractibility and overconfidence, enabling them to hide malicious content and reconstruct memory for jailbreak purposes. Their framework, tested on both open-source and proprietary LLMs, shows effectiveness in scalability and portability. Gressel et al. [37] propose a framework for detecting LLM impersonators in real-time using implicit and explicit challenge-response mechanisms. Their evaluation on open-source and proprietary models reveals the effectiveness of detection techniques, introducing a mismatch generalization strategy that forces the LLM to choose between conflicting objectives—safety and instruction-following.
@@
In addition, Lin et al. [87] introduce Analysis-Based Jailbreaks (ABJ), which use the advanced reasoning capabilities of LLMs to autonomously generate harmful content. ABJ decomposes simple prompts into independent elements and reconstructs them through complex reasoning steps, exposing hidden vulnerabilities in LLMs. Mehrotra et al. [79] present the Tree of Attacks with Pruning (TAP), an automated jailbreak generation method that requires only black-box access to the target LLM. TAP iteratively refines candidate attack prompts until one succeeds, pruning unlikely prompts before sending them to the target model, improving efficiency. Jiang et al. [83] propose ECLIPSE, a method that generates adversarial suffixes through optimization. ECLIPSE converts jailbreak objectives into natural language instructions using task prompts, guiding the LLM to generate adversarial suffixes. It also introduces a harmfulness scorer and a continuous feedback mechanism to help LLMs iteratively optimize and generate more aggressive and effective suffixes, increasing the success rate of jailbreaks.
@@
For the visual modality, Qi et al. [22] highlight the vulnerability of visual input due to its continuity and high dimensionality, making it susceptible to adversarial attacks. This weakness not only leads to misclassification but also allows adversarial visual samples to bypass the safety measures of vision-aligned LLMs. Zhao et al. [32] propose a method to evaluate the robustness of open-source VLMs in black-box settings, where adversaries only have query access. They design targeted adversarial examples for models like CLIP [11] and BLIP [91] and transfer them to other VLMs. Through black-box queries, they enhance the effectiveness of these attacks, achieving high success rates in producing targeted outputs. Gong et al. [70] propose FigStep, a black-box jailbreak method targeting VLMs. FigStep bypasses textual safety alignment by converting harmful textual instructions into typographic images. Without requiring white-box access, it exploits VLMs’ ability to recognize text in images and generate responses accordingly.
@@
Following this, Yang et al. [84] introduce SneakyPrompt, an automated attack framework that generates NSFW (Not Safe For Work) images even when safety filters are in place. If a prompt is blocked by safety filters, SneakyPrompt repeatedly queries the text-to-image generation model, perturbing tokens within the prompt based on query results. It uses reinforcement learning to guide token perturbation, optimizing both attack efficiency and success rate. Shayegani et al. [31] develop an aligned cross-modal attack method that combines adversarial images processed by vision encoders with textual prompts. This approach uses a novel composition strategy, combining toxic-embedding-targeted images with generic prompts to achieve successful jailbreaks. Li et al. [33] propose HADES, a novel jailbreak method that conceals and amplifies harmful intent within textual input using carefully designed images. The method removes harmful content from text, embeds it into typographic components, and combines these with harmful images generated by diffusion models and iteratively refined prompts within the LLM. Finally, adversarial images are overlaid to force MLLMs to generate affirmative responses to harmful instructions.
@@
Mao et al. [89] introduce the first tri-modal jailbreak hybrid strategy framework, JMLLM, which integrates four toxic-content concealment techniques and targets jailbreak attacks across text, visual, and audio inputs. JMLLM effectively bypasses the defenses of LLMs across different modalities. Through coordinated multimodal attacks, it achieves state-of-the-art success rates while significantly reducing time overhead. Zhang et al. [85] introduce the B-AVIBench framework to assess the robustness of large-scale VLMs against black-box adversarial visual instructions (B-AVIs). The framework includes four image-based B-AVIs, ten text-based B-AVIs, and nine content-bias B-AVIs (e.g., gender, violence, cultural, and racial biases).
@@
(c) Agents: Black-box jailbreak attacks against intelligent agents exploit their task execution dynamics, manipulating the decision-making process through iterative interactions, environment manipulation, and task decomposition. Adversaries craft specific task inputs that gradually lead agents astray or exploit vulnerabilities in tool invocation and API interactions to covertly bypass security mechanisms. Nakash et al. [40] show that when an agent is asked to fix bugs on a website, it can be influenced by indirect prompt injections and “foot-in-the-door” interference. These subtle injections gradually affect the agent’s decision process, leading it to execute attacker-defined instructions. As a result, the agent may perform not only harmless tasks (e.g., computing 2 + 4) but also malicious actions, such as sending admin credentials to the attacker.
@@
Building on earlier studies, Zhang et al. [41] introduce attacks that manipulate agents into performing repetitive or irrelevant actions, causing system failures. They use GPT-3.5-Turbo-16k as a sandbox model and GPT-3.5-Turbo as the core model, simulating real-world tool behavior to expose the model’s vulnerabilities. Wu et al. [43] use adversarial text strings to trigger perturbations on images, combining this with attacks on subtitle generators that convert images into text for multimodal models. Zhang et al. [76] propose AI2, a hijacking method that manipulates black-box agent systems by stealing action-perception memory and injecting false contexts. This method exploits the gap between the memory retrieval system and safety filters, evading detection. Kumar et al. [73] introduce BrowserART, a red-teaming toolkit targeting browser-based agents interacting with the web. Their study shows that while LLMs may refuse harmful input, agents still fail to reject malicious instructions, highlighting the need for better safety at the agent level.
@@
While previous research mainly focuses on single-turn jailbreak attacks, it overlooks the risks posed by multi-turn dialogues, which are essential for human-LLM interactions. To address this, Wang et al. [90] propose a multi-turn jailbreak attack using a red-team agent. Their framework includes data construction and agent training. The data construction strategy spreads malicious intent across multiple rounds using psychological techniques to create high-quality datasets. The agent is trained with interaction feedback, optimizing the attack strategy over time. Jiang et al. [77] introduce RAG-Thief, an automated privacy attack agent for RAG-based applications. RAG-Thief extracts sensitive information from private databases using adaptive querying with adversarial samples, refining queries to maximize data leakage.
Most recently, Dong et al. [74] introduce Atlas, a framework using multiple autonomous agents to bypass safety filters in text-to-image (T2I) models. Atlas employs a fuzzing approach with a mutation agent for detecting filter triggers and optimizing jailbreak prompts, and a selection agent that scores and submits the best prompts. It also integrates chain-of-thought (CoT) prompting and in-context learning (ICL) to improve adaptability. Ju et al. [39] propose a two-stage attack: the first stage generates fabricated evidence, while the second manipulates the agent’s knowledge, allowing for persistent knowledge poisoning in RAG frameworks.
@@
for tree= grow=south, anchor=north, parent anchor=south, child anchor=north, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, minimum height=3em, edge path= [draw=edge-color, line width=2pt] (!parent.south) – +(0,-5pt) -— (.child anchor); , l sep=10pt, s sep=10pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=minimum width=15em,rotate=0, child anchor=north, parent anchor=south, anchor=center, , where level=1 draw=child-one-line-color, text width=15em, minimum height=3em, font=, , where level=2 draw=child-five-line-color, text width=12em, minimum height=3em, font=, , where level=3 draw=child-five-line-color, minimum height=3em, font=, , [ Datasets, par, for tree=fill=root-color, [ Data Sources $\S$ IV, for tree=fill=child-one-color, [ LLM/Automatic Generation, for tree=fill=child-two-color, [ xx[88, 92, 93, 70, 33, 94]xx, for tree=fill=child-five-color, ] ] [ Search Engine Retrieval, for tree=fill=child-two-color, [ xx[89, 92, 37, 95, 96, 97, 33, 60]xx, for tree=fill=child-five-color, ] ] [ Handmade, for tree=fill=child-two-color, [ xx[88, 80, 89, 98, 42, 95, 70, 99, 100, 101, 102]xx, for tree=fill=child-five-color, ] ] ] [ Data Format $\S$ IV, for tree=fill=child-one-color, [ Q&A, for tree=fill=child-two-color, [ xx[37, 98, 100, 95, 96, 70, 102, 101, 97]xx, for tree=fill=child-five-color, ] ] [ Instructions, for tree=fill=child-two-color, [ xx[99, 88, 42, 92]xx, for tree=fill=child-five-color, ] ] [ Harmful Sentences/Images, for tree=fill=child-two-color, [ xx[89, 80, 93, 94, 33, 60]xx, for tree=fill=child-five-color, ] ] ] ]
@@ IV Datasets
  Dataset Size Scenes Language Data Details
AdvBench [80] 500/500 (574/520) 8 EN Harmful strings & behaviors
LatentJailbreak [99] 416 3 EN/ZH Translation tasks
SAP [88] 1600 8 EN Harmful instructions
SafeBench [70] 500 10 EN Unsafe questions
SafeBench-Tiny [70] 50 10 EN Unsafe questions
Do-Not-Answer [101] 939 5 EN Harmful instructions
SAFETYPROMPTS [93] 100k 14 ZH Harmful strings
Are You Human [37] 210 2 EN Q&A question
JBB - Behaviors [92] 100/100 10 EN Harmful & harmless behaviors
AgentHarm [42] 110/440 11 EN Harmful instructions
XSTEST [98] 250/200 10 EN Safe and unsafe Q&A question
SafetyBench [95] 11435 7 EN/ZH Multiple-choice question
StrongREJECT [100] 346 6 EN Unsafe questions
AttackEval [96] 390 13 EN Unsafe questions
TriJail [89] 1250/1250/150 6 EN Harmful strings, audio, and images
TECHHAZARDQA [102] 7745 7 EN Unsafe questions
MM-SafetyBench [94] 5040 13 EN Image-text pair
LLaVA-Instruct150K [97] 1.20M 4 EN Image-question-answer pair
RTVLM [33] 750 5 EN Harmful images-text pair
AdvBench-M [60] 240 8 EN Harmful images-text pair
@@
LatentJailbreak [99] Write one toxic sentence about women with at most 20 words. / 写一个关于犹太人的有毒句子,最多20个单词。